# Depth-Structured Music Recurrence: Budgeted Recurrent Attention for Full-Piece Symbolic Music Modeling

### v1

@@ VI-B Deterministic deep processing cannot improve the Bayes-optimal log-loss
Consistent with this analysis, our experiments show that the choice of which layers to retain in DSMR (shallower versus deeper) does not lead to a substantial difference. Within the two-scale DSMR group, retaining a shallower layer is slightly better than retaining a deeper one. This suggests that carrying “more contextual information” via deeper processing does not inherently yield lower NLL, and helps explain why the two-scale DSMR variant that retains only shallow layers for long context performs best in our setting.
VI-C State Staleness and Optimization Stability
@@
In our stateful recurrent training, later segments condition on cached cross-segment states computed under slightly older parameters, so the recurrent context is stale relative to the parameters used for the current update. Empirically, extending the recurrence horizon beyond $30$ segments, which is substantially longer than the effective history used in the original Transformer-XL experiments [12], does not degrade convergence in the DSMR setting.
@@
This behavior can be understood through the notion of bounded staleness, as studied in pipelined and asynchronous optimization, where computations may involve delayed parameter versions while remaining stable under controlled updates [25, 23, 37]. In our setting, early training typically spans on the order of $300\text{k}$ optimization steps. Under Adam-style optimization, the parameter change at each update is small, so the discrepancy between stale and freshly recomputed recurrent states behaves as a first-order perturbation. Rather than introducing systematic bias, this discrepancy acts as mild structured noise whose magnitude is bounded by the update scale.
This stability provides a practical foundation for training recurrent models end to end on full-piece music. Even when later segments depend on states produced many updates earlier, the induced inconsistency remains controlled, enabling long-horizon recurrence without compromising optimization dynamics.
VI-D Practical implications for long-context music modeling
Our experiments are conducted on MAESTRO, where long-range musical coherence often requires context spanning many bars or an entire performance. From a practical perspective, the combination of full-piece end-to-end training, stateful recurrence and a fixed memory-depth budget provides an effective recipe to train long-context models under constrained compute.
@@ VI-D Practical implications for long-context music modeling
This is particularly relevant to music applications with tight latency or hardware limits (e.g., interactive composition tools and rehearsal/performance-time assistance), where reducing both time-axis cost and the cost of attending over long cross-segment histories at many layers is critical. A representative scenario is real-time improvisational interaction with a human musician: the system continually uses the musician’s just-played phrase as a primer and generates a responsive continuation for turn-taking or call-and-response, a workflow that is especially common in jazz performance. Such settings often require the model to run on portable, resource-constrained devices, where our approach is well suited by design.
VI-E Limitations
Our findings are bounded by the regimes explored: validation PPL is teacher-forced and may not fully reflect long-range structural generation quality; the conclusions may depend on the model scale, memory/segment settings, and task/dataset may affect substitutability; and we do not exhaust the design space of memory-retention or gating, evaluate only a limited set of heuristic horizon combinations and selecting the best configuration within the explored set.
VII Conclusion
We presented Depth-Structured Music Recurrence (DSMR) for full-piece symbolic music modeling, extending context beyond fixed-length excerpts through segment-level stateful recurrence with detached cross-segment KV states, and allocating recurrent memory across depth via a layer-wise horizon schedule. On MAESTRO, DSMR shows that combining full-piece end-to-end training with stateful recurrence, and explicitly redistributing recurrent history across depth via a layer-wise memory-horizon schedule, yields a strong quality–efficiency tradeoff for long-context symbolic music modeling. A simple two-scale DSMR schedule, which assigns long horizons to lower layers and a uniform short horizon elsewhere, is the best-performing configuration within the heuristic schedules explored, achieving slightly lower perplexity than a full-memory reference while using much less memory and enabling faster training.
From an information perspective, longer context provides a richer information structure for prediction and cannot worsen the optimal achievable log-loss when the model family can ignore irrelevant tokens. Moreover, when a “deep” context is obtained only by deterministic processing of the full context without injecting extra information, information-theoretic limits imply that such deep processing can at best match, but cannot improve upon, the Bayes-optimal log-loss achievable when the full context is made available already at a shallow stage. This supports the practical DSMR design choice of retaining full-length context via relatively shallow processing, while allocating shorter receptive fields to higher layers under a fixed recurrent-state budget.
@@ VII Conclusion
Future work includes evaluating DSMR under more extreme capacity bottlenecks, and assessing its generality beyond symbolic music by testing DSMR on other modalities and long-context domains where stateful recurrent attention is relevant.
@@ Appendix A Stateful Recurrence and Long-Context Attention Approximations
A-A Transformer-XL style stateful recurrence
@@
Transformer-XL maintains a per-layer KV memory from previous segments [12] (Fig. 9). Let $M_{t-1}^{(\ell)}\in\mathbb{R}^{m\times d}$ be the memory used at layer $\ell$ for segment $t$ (memory length $m$). When computing self-attention in layer $\ell$, it concatenates historical memory and current hidden states:
$\widetilde{H}_{t}^{(\ell-1)}=\bigl[\mathrm{SG}(M_{t-1}^{(\ell-1)});\ H_{t}^{(\ell-1)}\bigr]\in\mathbb{R}^{(m+s)\times d},$ (12)
@@
$H_{t}^{(\ell)}=\mathrm{Block}^{(\ell)}\!\left(H_{t}^{(\ell-1)},\ \widetilde{H}_{t}^{(\ell-1)}\right),$ (13)
where keys/values come from $\widetilde{H}_{t}^{(\ell-1)}$ and queries come from $H_{t}^{(\ell-1)}$ (Fig. 9).
Memory is updated by keeping the most recent $m$ token states:
$M_{t}^{(\ell)}=\mathrm{Last}_{m}\!\left(\bigl[M_{t-1}^{(\ell)};\ H_{t}^{(\ell)}\bigr]\right).$ (14)
Figure 9: Transformer-XL style stateful recurrence with stop-gradient KV memory. At each layer, keys/values (green) concatenate cached history with the current segment, while queries (purple) are computed for the current segment only.
A-B How recurrence approximates segment-wise full attention
@@
$\mathrm{Attn}\!\Bigl(Q_{t},\ [K_{<t};K_{t}],\ [V_{<t};V_{t}]\Bigr).$ (15)
Figure 10: Segment-wise full attention view. For queries from the current segment, full attention uses keys/values from both the historical prefix and the current segment (Eq. 15). See Fig. 9 for the stateful recurrence implementation that approximates this form in the forward pass when memory is untruncated.
Stateful recurrence approximates Eq. (15) by caching historical activations so the model need not recompute the entire prefix at each step. With untruncated memory, $M_{t-1}^{(\ell-1)}$ supplies historical representations as keys/values together with the current segment (Eqs. (12)–(13)), yielding the same forward-pass form as Eq. (15). As illustrated in Fig. 10, the concatenated KV source $\bigl[M_{t-1}^{(\ell-1)};H_{t}^{(\ell-1)}\bigr]$ plays the same role as $\bigl[K_{<t};K_{t}\bigr]$ (and likewise for values), matching Eq. (15) in the forward pass. Training, however, differs from true full attention due to stop-gradient through memory and statefulness: historical states consumed at step $t$ are cached activations computed using parameters from earlier optimization steps and treated as constants, which shapes how long-range capability is learned under fixed budgets.
A-C Perceiver AR as a stateless long-KV / short-Q analogue
Perceiver AR is not a stateful recurrent Transformer, but its first attention stage exhibits a similar long-KV / short-Q pattern to Transformer-XL-style recurrence [19]. Specifically, Perceiver AR uses a causally masked input cross-attention that maps a long input sequence into a much shorter latent array, typically aligning one latent with each of the final $N$ input positions. Let $E_{1:T}\in\mathbb{R}^{T\times d}$ be input embeddings and $Z\in\mathbb{R}^{N\times d}$ be the latent queries. The cross-attention can be written abstractly as
$\mathrm{Attn}\!\bigl(Q_{\mathrm{lat}},\,K_{\mathrm{in}},\,V_{\mathrm{in}}\bigr),$ (16)
@@ A-C Perceiver AR as a stateless long-KV / short-Q analogue
The key difference is statefulness. In Transformer-XL-style recurrence, the long-range KV comes from cached hidden states carried across segments (and detached via stop-gradient), whereas in Perceiver AR the long-range KV is formed by projecting the current input sequence embeddings at each forward pass (stateless), with causality enforced by masking. Figure 11 illustrates the analogy.
Figure 11: Perceiver AR as long-KV / short-Q cross-attention [19]. A causally masked cross-attention maps a long input (projected embeddings as keys/values) to a short latent array (queries), after which a deep stack of self-attention operates only on the short latents. This is analogous to exposing long-range KV to an early stage, but differs from recurrent Transformers in that the KV is stateless (recomputed each pass) rather than cached recurrent state.

### v2

@@ 6.1 Information-theoretic perspective on context depth
Consistent with this analysis, our experiments show that retaining long context at a shallower layer slightly outperforms retaining it at a deeper one. This helps explain why two-scale DSMR, which concentrates long horizons in lower layers, is the best-performing schedule in our setting.
6.2 State staleness and optimization stability
@@
Because DSMR uses truncated backpropagation with a stop-gradient across segments (Eq. (4)), later segments condition on cached states produced under slightly older parameters. In our setting, each piece spans at most $T\leq\lceil 32768/s\rceil=32$ recurrent steps (where s=1024s{=}1024 is the segment length), while training runs for roughly 33M optimizer steps; the step-to-step parameter mismatch over the cached history is therefore small. Empirically, extending the recurrence horizon beyond $30$ segments—substantially longer than the effective history in the original Transformer-XL experiments Dai et al. [2019]—does not degrade convergence.
@@
This behavior is consistent with bounded staleness in pipelined and asynchronous optimization, where delayed parameter versions remain stable under controlled updates Narayanan et al. [2019], Kosson et al. [2021], Zhang et al. [2016]. Under Adam-style optimization, the per-step parameter change is small, so the discrepancy between stale and freshly recomputed states behaves as a first-order perturbation rather than systematic bias. This stability provides a practical foundation for end-to-end full-piece training with long-horizon recurrence.
6.3 Practical implications for long-context music modeling
@@
DSMR is designed around creator-centric, resource-limited workflows: depth-wise horizon budgeting keeps full-piece end-to-end training viable on consumer-grade hardware. This is relevant to music applications with tight hardware limits, such as interactive composition tools and rehearsal-time assistance. A representative scenario is real-time improvisational interaction, where the system uses a musician’s just-played phrase as a primer and generates a responsive continuation for turn-taking or call-and-response—a workflow common in jazz performance. Such settings often require models to run on portable devices, where the memory and throughput savings demonstrated by DSMR are directly beneficial. We note, however, that our evaluation uses offline teacher-forced perplexity; validating DSMR in real-time generation settings remains future work.
6.4 Limitations
Our findings are bounded by the regimes explored. (i) Validation PPL is teacher-forced and may not fully reflect long-range structural generation quality; no listening test or perceptual evaluation is conducted. (ii) Experiments use a single dataset (MAESTRO) and a single instrument (piano); conclusions about layer substitutability and depth placement may not generalize to other musical genres, ensemble textures, or non-music domains. (iii) We evaluate only one model scale (L=18L{=}18, d=1024d{=}1024); the optimal horizon schedule may differ at larger or smaller capacities. (iv) No comparison is made against non-recurrent long-context methods (e.g., sparse attention trained on full pieces).
7 Conclusion
We presented Depth-Structured Music Recurrence (DSMR), a training-time design that enables full-piece symbolic music modeling by streaming complete compositions left-to-right with stateful recurrent attention under a fixed layer-wise horizon budget. On MAESTRO, a two-scale DSMR schedule—long horizons in lower layers, a uniform short horizon elsewhere—matches the full-memory reference in perplexity (5.96 vs. 5.98) while using 59% less GPU memory and achieving 36% higher throughput. Variant analyses reveal strong layer substitutability under binary-horizon schedules and show that gate-derived layer rankings do not transfer reliably under structural intervention. Information-theoretically, these findings align with the principle that longer context cannot worsen optimal log-loss and that deterministic deep processing cannot surpass the Bayes limit, supporting the design choice of concentrating long-range access at shallow layers.
@@ 7 Conclusion
Future work includes evaluating DSMR under more extreme capacity bottlenecks and at larger model scales, assessing perceptual generation quality through listening tests, and testing DSMR on other modalities and long-context domains where stateful recurrent attention is relevant.
@@ Appendix A Stateful Recurrence and Long-Context Attention Approximations
A.1 Transformer-XL style stateful recurrence
@@
Transformer-XL maintains a per-layer KV memory from previous segments Dai et al. [2019] (Fig. A.1). Let $M_{t-1}^{(\ell)}\in\mathbb{R}^{m\times d}$ be the memory used at layer $\ell$ for segment $t$ (memory length $m$). When computing self-attention in layer $\ell$, it concatenates historical memory and current hidden states:
$\widetilde{H}_{t}^{(\ell-1)}=\bigl[\mathrm{SG}(M_{t-1}^{(\ell-1)});\ H_{t}^{(\ell-1)}\bigr]\in\mathbb{R}^{(m+s)\times d},$ (A.1)
@@
$H_{t}^{(\ell)}=\mathrm{Block}^{(\ell)}\!\left(H_{t}^{(\ell-1)},\ \widetilde{H}_{t}^{(\ell-1)}\right),$ (A.2)
where keys/values come from $\widetilde{H}_{t}^{(\ell-1)}$ and queries come from $H_{t}^{(\ell-1)}$ (Fig. A.1).
Memory is updated by keeping the most recent $m$ token states (Eq. (1)).
Figure A.1: Transformer-XL style stateful recurrence with stop-gradient KV memory. At each layer, keys/values (green) concatenate cached history with the current segment, while queries (purple) are computed for the current segment only.
A.2 How recurrence approximates segment-wise full attention
@@
$\mathrm{Attn}\!\Bigl(Q_{t},\ [K_{<t};K_{t}],\ [V_{<t};V_{t}]\Bigr).$ (A.3)
Figure A.2: Segment-wise full attention view. For queries from the current segment, full attention uses keys/values from both the historical prefix and the current segment (Eq. (A.3)). See Fig. A.1 for the stateful recurrence implementation that approximates this form in the forward pass when memory is untruncated.
Stateful recurrence approximates Eq. (A.3) by caching historical activations so the model need not recompute the entire prefix at each step. With untruncated memory, $M_{t-1}^{(\ell-1)}$ supplies historical representations as keys/values together with the current segment (Eqs. (A.1)–(A.2)), yielding the same forward-pass form as Eq. (A.3). As illustrated in Fig. A.2, the concatenated KV source $\bigl[M_{t-1}^{(\ell-1)};H_{t}^{(\ell-1)}\bigr]$ plays the same role as $\bigl[K_{<t};K_{t}\bigr]$ (and likewise for values), matching Eq. (A.3) in the forward pass. Training, however, differs from true full attention due to stop-gradient through memory and statefulness: historical states consumed at step $t$ are cached activations computed using parameters from earlier optimization steps and treated as constants, which shapes how long-range capability is learned under fixed budgets.
A.3 Perceiver AR as a stateless long-KV / short-Q analogue
Perceiver AR is not a stateful recurrent Transformer, but its first attention stage exhibits a similar long-KV / short-Q pattern to Transformer-XL-style recurrence Hawthorne et al. [2022]. Specifically, Perceiver AR uses a causally masked input cross-attention that maps a long input sequence into a much shorter latent array, typically aligning one latent with each of the final $N$ input positions. Let $E_{1:T}\in\mathbb{R}^{T\times d}$ be input embeddings and $Z\in\mathbb{R}^{N\times d}$ be the latent queries. The cross-attention can be written abstractly as
$\mathrm{Attn}\!\bigl(Q_{\mathrm{lat}},\,K_{\mathrm{in}},\,V_{\mathrm{in}}\bigr),$ (A.4)
@@ A.3 Perceiver AR as a stateless long-KV / short-Q analogue
The key difference is statefulness. In Transformer-XL-style recurrence, the long-range KV comes from cached hidden states carried across segments (and detached via stop-gradient), whereas in Perceiver AR the long-range KV is formed by projecting the current input sequence embeddings at each forward pass (stateless), with causality enforced by masking. Figure A.3 illustrates the analogy.
Figure A.3: Perceiver AR as long-KV / short-Q cross-attention Hawthorne et al. [2022]. A causally masked cross-attention maps a long input (projected embeddings as keys/values) to a short latent array (queries), after which a deep stack of self-attention operates only on the short latents. This is analogous to exposing long-range KV to an early stage, but differs from recurrent Transformers in that the KV is stateless (recomputed each pass) rather than cached recurrent state.
@@
Acknowledgements
AI-Assisted Language Editing. During the preparation of this work the authors used ChatGPT in order to improve English grammar and readability. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.